Arizona LBTI Commissioning - 17 October 2010 UT

This wiki page includes John's notes on the center bent Gregorian focal station commissioning aspects of LBTI Commissioning. This represents only a small fraction of the total activity.

Observers: JHill, PHinz, Vidhya, Bill, Vanessa
Telescope Operators: SAllanson, GBechetti
Instrument Support: JMorris
Telescope Support: MMidkiff, BCurtis

Summary

Variable clouds and variable seeing, but we've had fringes on the detector nearly all night long. We took a brief look at 4.8 micron fringes and found the easily. Collected various calibration data with MIRAC and worked on various software issues. Sorted out some details about sign conventions for repointing and overlapping the images. Had some failure? of the LBTI piston adjusting stage in the last hour of the night.

Preparations

Details

Open

03:10 Open as clouds start to clear. T=5.5 degC D=1.4 degC Wind light.

03:24 LBTI TRACK Preset to Beta Peg

Had problems with SX M3 moving out of collimation position again, but it only seems to happen once per night ??

Found both stars at IE=-44 CA=+38. As before, we are using DX coma-free repointing to overlap the images. DX M1M2 is Rx=-3 Ry=+39

From eyeball active optics, SX needed nothing but -0.12 mm of focus, but DX needed Z5, Z6, Z7 and Z8 (and still looks worse than SX). Seeing is much better tonight, so both images show a nice Airy core at 10 microns.

We now have images on both acquisition cameras although the centering or mis-centering relative to MIRAC is not completely understood.

Search for fringes

04:05 We got the first fringes well before midnight tonight. When the DX mirror collimates, we lose the image. This is because the DX Collimation Model has different pointing terms at high elevation. In principle the accumulated Rx Ry terms at different elevations will let us correct the DX collimation model, but see the summary at the bottom of this page. There turns out not to be a very strong trend with elevation.

04:15 EL=83 IE=-46 CA=+38 DX Rx=0 Ry=34

04:30 EL=84 IE=-43 CA=+36 DX Rx=0 Ry=34

Phil recenters the acquisition stages so everybody is lined up. We have photons on both acquisition cameras and both pyramids.

04:55 Geno reports seeing is 1 arcsec on the DIMM. It's even better at 10 microns.

05:10 EL=82 IE=-49 CA=38 DX Rx=0 Ry=+43

05:45 EL=74 IE=-57 CA=+36 DX Rx=0 Ry=+43

05:49 Vidhya sends a 0,0 guide command, and that moves the stars away to an unknown location.

05:52 Resend preset

05:58 Autoguide 0,0 correctly does nothing, but a second command doesn't make it through. Vidhya restarted her server, and then the commands got through. But they seem to be acting in RADEC rather than ALTAZ.

06:02 Commands that Phil issues from the LBTI GUI don't appear at the telescope for a couple minutes.

06:20 Clouds and humidity from the SE are threatening.

06:37 EL=64 IE=-65 CA=+42 DX Rx=-4 Ry=43

Phil finds the fringes down here only a fraction of a millimeter from where they were 20 degrees higher in elevation. If we can figure out the sign, we could include that change in the collimation model.

07:01 EL=59 IE=-65 CA=+42 DX Rx=-5 Ry=+45

Now taking some 4 millisec integrations at 100 Hz. But seeing may be worse.

07:06 Seeing on DIMM at 1 arcsec.

We are doing "Azimuth Nodding" (term used loosely) for background measurements by offsetting 20 arcsec in CA (or sometimes IE when I forget).

07:54 Eyeball collimation at EL=48. Mostly needed more negative Z6 on DX.

07:59 Guide commands didn't produce the expected symmetric result. The star didn't come back after an out/back motion.

08:00 resend preset found star at EL=46 IE=-69 CA=+35 noticeably different than previous IE/CA on same star. We think this was because Phil had been essentially guiding with offset commands, so John hadn't made any IE/CA changes for a while. Maybe LBTI should guide with IE and CA all the time?

08:19 EL=43 IE=-71 CA=+38 DX Rx=-4 Ry=+45

08:20 Vidhya's first moveFocus command arrives.

08:32 Preset to calibrator star Alpha Tau, EL=64 IE=-62 CA=38 DX Rx=0 Ry=52
SX Zglob=0.095 DX Zglob=-1.05

Note that the new control room walls don't make popping noises during azimuth slews.

08:50 Sky is mostly clear, but seeing is worse.

Mirror Ventilation is not getting enough coolth, so Steve turned it off before the mirrors heat up to much. See Issue 2935.

08:31 SX Zglob=0.0 DX Zglob=0.0 DX Rx=0 Ry=48

09:19 Fringes at 4.8 microns.

09:28 EL=71 IE=-62 CA=43

Guiding Tests

09:40 Phil sends +10 -10 guide commands, and star doesn't come back. Did the offset commands wipe out the guide origin?

09:45 EL=72 IE=-55 CA=36

10:00 Seeing is now much worse, so we can't see 10 micron fringes anymore.

10:24 With refocussing and slightly better seeing, we have fringes again.

EL=73 IE=-58 CA=36 DX Rx=0 Ry=45

10:38 EL=72 IE=-60 CA=34 DX Rx=0 Ry=44 Fringes are still there without any manual pathlength adjustment for an hour. I am keeping the images overlapped manually, since Vidhya is still working on the guiding software and this is monocular PCS.

Repointing Tests

10:50 We've lost connection with abell.as.arizona.edu, but we can still talk to wiki.lbto.arizona.edu

10:55 EL=71.2 IE=-63 CA=33 DX Rx=0 Ry=+43

See the notes on which pointing changes move the star in which direction in Com20101015. See also Issue 2934 about a future sign change in DX Ry.

12:02 SX Zglob=0.16 DX Zglob=-0.9

Trying for Binary Star Fringes

12:02 Preset to Alpha Gem EL=77 IE=-64 CA=38

12:22 EL=79 IE=-62 CA=+37 DX Rx=-1 Ry=+40

12:47 restart motor controller on piston mirror, but it didn't work. So no fringe data on Alpha Gem tonight.

12:55 Close for impending dawn and busted piston motor controller. T=4.5 degC D=0.4 degC Sky is Clear. Wind is 5 m/sec from East.

Trend of copointing with Elevation?

Here's the reordered list of copointing parameters from the notes above. It turns out that the trends with elevation are not very strong.
04:30 EL=84 IE=-43 CA=+36 DX Rx=0 Ry=34

05:10 EL=82 IE=-49 CA=+38 DX Rx=0 Ry=+43

12:22 EL=79 IE=-62 CA=+37 DX Rx=-1 Ry=+40

05:45 EL=74 IE=-57 CA=+36 DX Rx=0 Ry=+43

10:26 EL=73 IE=-58 CA=+36 DX Rx=0 Ry=+45

10:38 EL=72 IE=-60 CA=+34 DX Rx=0 Ry=+44

09:45 EL=72 IE=-55 CA=+36

10:55 EL=71 IE=-63 CA=+33 DX Rx=0 Ry=+43

09:28 EL=71 IE=-62 CA=+43

06:37 EL=64 IE=-65 CA=+42 DX Rx=-4 Ry=+43

08:32 EL=64 IE=-62 CA=+38 DX Rx=0 Ry=+52

07:01 EL=59 IE=-65 CA=+42 DX Rx=-5 Ry=+45

08:00 EL=46 IE=-69 CA=+35

08:19 EL=43 IE=-71 CA=+38 DX Rx=-4 Ry=+45

-- JohnHill - 17 Oct 2010
Topic attachments
I Attachment Action Size Date Who Comment
pmcc.dx.thermal.20101017.jpgjpg pmcc.dx.thermal.20101017.jpg manage 548 K 18 Oct 2010 - 10:56 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20101017.jpgjpg pmcc.sx.thermal.20101017.jpg manage 523 K 18 Oct 2010 - 10:57 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r8 - 18 Oct 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback